I need help writing the slope-intercept equation for the line that passes through (-3, -15) and is perpendicular to -6x + 8y = 3. Thank you.
Answer by mananth(16946)

(-3, -15) and is perpendicular to -6x + 8y = 3. Thank you.
-6x+8y = 3
Find the slope of this line
8y=6x-3
Divide by 8
y= 3/4x-0.38
Compare this equation with y=mx+b
slope m =0.75
The slope of a line perpendicular to the above line will be the negative reciprocal = -1.33
(m1*m2=-1) m=
-1.33 ,point( -3,-15)
Find b by plugging the values of m & the point in
y=mx+b
-15=4.00 +b
b=-19
m=-1.33
The required equation is y=-1.33x -19
m.ananth@hotmail.ca